Calculate the number of moles of Na+ ions in 25 g of sodium carbonate, Na2 CO3.


1
Expert's answer
2021-05-08T23:33:51-0400

Solution:

Na2CO3 → 2Na+ + CO32-

According to the equation above: n(Na2CO3) = n(Na+)/2


The molar mass of Na2CO3 is 105.9888 g/mol.

Hence,

(25 g Na2CO3) × (1 mol Na2CO3 / 105.9888 g Na2CO3) = 0.23587 mol Na2CO3

n(Na2CO3) = 0.23587 mol


n(Na+) = 2 × n(Na2CO3) = 2 × 0.23587 mol = 0.47174 mol Na+ ions = 0.472 mol Na+ ions

Moles of Na+ ions = 0.472 mol


Answer: 0.472 moles of Na+ ions in 25 g of sodium carbonate (Na2CO3).